OpenStudy (anonymous):

4 l 2y - 7 l + 5 = 9

OpenStudy (lgbasallote):

\[4 | \; 2y - 7 \; | + 5 = 9\] subtract 5 from both sides \[4 | \; 2y - 7 \; | = 4\] divide both sides by 4 \[| \; 2y - 7 \; | = 1\] now solve for y in: \[2y - 7 = 1\] AND \[2y - 7 = -1\] does that help?

OpenStudy (anonymous):

So the answer is 1 and also -1?

OpenStudy (lgbasallote):

no. solve for y

OpenStudy (lgbasallote):

what's y in 2y - 7 = 1?

OpenStudy (anonymous):

Oh sorry okay so is the answer 7/2 and 3?

OpenStudy (anonymous):

4

OpenStudy (lgbasallote):

no... 2y - 7 = 1 solve for y what do you get?

OpenStudy (lgbasallote):

yes 4

OpenStudy (lgbasallote):

now.. 2y - 7 = -1 solve for y what do you get?

OpenStudy (anonymous):

3

OpenStudy (lgbasallote):

so y = 4 and y = 3

OpenStudy (lgbasallote):

^that's the answer

OpenStudy (anonymous):

Oh okay thank you!

OpenStudy (lgbasallote):

welcome